OpenStudy (anonymous):

Which of the following is the solution to the equation 8n = 125(n + 4) ?

OpenStudy (anonymous):

first thing you want to do is look to the right of the "=" sign and multiply 25 by what is in the bracket 8n = 125n + 500

OpenStudy (anonymous):

then you add the like terms right?

OpenStudy (anonymous):

you have to have the "n" on one side of the "=" sign and the numbers on the other

OpenStudy (anonymous):

i messed the question up in front of the 8n it should be sqaure root of 5^8n= 125^(n+4)

OpenStudy (anonymous):

8n - 125n = 500 - 117 n = 500 n = -500/117

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

I think you should repost your question with the corrected problem...Im not 100% sure on how to do the corrected problem

OpenStudy (anonymous):

maybe once I see the corrected problem again, more clearly, I might be able to solve it

OpenStudy (anonymous):

never mind i got it thanks for the help